Operation Summary for February 14, 2011 - February 16, 2011 |
|
Monday Feb. 14 The day shift began with Operations monitoring the Tevatron (TEV) store 8498 (156.27E30) and the Antiproton Source (Pbar) stack (9.5E10), with no Recycler stash and with the Main Injector (MI) off due to a MI ground fault indication, with the Linac and Booster operating normally, with the NuMI experiment off due to MI problems, with the MiniBooNE experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Interlock personnel reported tracing a MI safety system problem to a bad card, which they replaced. MI resumed sending beam shortly after 10 AM. Operations reported that a ramp card issue held off stashing; Controls personnel believed that this was left over from the power glitch during the weekend. NuMI personnel conducted target scans and MI personnel conducted gap clearing kicker studies. During the evening shift, Operations terminated store 8498 at 10:11 PM. Unfortunately, the TEV quenched shortly thereafter. Neither operators nor TEV personnel could find a cause for the quench. The cryo system recovery continued through to the end of shift. Operations reported that the Pelletron had tripped off multiple times. Electron Cooling personnel came in to investigate; they decided the problem was inside the Pelletron and an access would be needed on Tuesday morning. At 7:09 PM, all the MI kickers tripped off and Booster extraction positions jumped considerably. Booster personnel came in and determined there was a bad power supply for a crate in the Booster low-level RF system. They began changing the power supply at the end of the shift. Operations reported that a Linac Klystron RF station had tripped off. Tuesday Feb. 15 The midnight shift began with the TEV recovering from a quench, with Operations monitoring the stash (198.6E10) and the Pbar stack (88.0E10), with NuMI and MiniBooNE off due to Booster RF problems. During the midnight shift, Operators prepared CDF for an hour access at the beginning of shift. Booster personnel had completed their crate power supply repairs by 12:42 AM. RF personnel completed repairs on a TEV RF station power amplifier at 1 AM. Cryo system personnel at CHL reported at 3 AM that the cryo system had recovered and the TEV was cold. At 2:53 AM, the Booster chopper power supply tripped off. Booster personnel came in and determined that they needed an access; they replaced a faulty termination. At 3:27 AM, a "network storm" hit. Controls personnel came in; they located and isolated the source and had the network back up by 5:10 AM. Unfortunately, the Booster gradient magnet power supply crate did not come back up. Booster personnel came in to help with its recovery. Operations had beam restored to the complex at 6:10 AM. Then the TEV lost the ramp permit on high-energy ramps. Cryo system personnel had operators adjust several valves at TEV sector C4 to increase the inlet pressure to the wet engine. During the day shift, Operations put the TEV into shot setup at 8:42 AM and established store 8502 at 10:19 AM with an AIL of 135.31E30. Preaccelerator personnel switched from the H- Source to the I- Source during the morning and tuned the new source for the rest of the shift. Operations reported several down times caused by Linac RF stations, a MI damper reboot, and problems with Booster quadrupole PLCs. During the evening shift, Operations terminated store 8502 at 10:11 PM, put the TEV into shot setup, and established store 8503 at 11:19 PM with an AIL of 116.34E30. Switchyard personnel conducted dump studies. Operations reported minor downtimes due to the Linac RF system. Wednesday Feb. 16 The midnight shift began with Operations monitoring store 8503 (93.97E30) and the stash (163.5E10), with Pbar stacking (39.2E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 8503 at 7:03 AM. Operations prepared D0 for an expected eight-hour access. Repairs on the Pelletron continued throughout the shift. Operators reported large Booster extraction position changes towards the end of shift; Booster personnel investigated. Operators reported minor down times due to Booster and Linac RF station problems. The Plans for Wednesday The Run Coordinator will allow some necessary machine repairs during the D0 access. |
|
More Information Comments and Suggestions |
| last modified 02/16/2011 email Fermilab |
| Security, Privacy, Legal | |